OpenStudy (anonymous):

Which property is shown? (8 • 5) • 11 = 8 • (5 • 11) Simplify the expression. 7d + 12– 4d – 3

OpenStudy (anonymous):

thx

OpenStudy (hba):

7d+12-4d-3=0 lets solve it :)

OpenStudy (hba):

@-Lexy- Start working

OpenStudy (anonymous):

cant xD im terrible at this

OpenStudy (hba):

Let me show you an example 3c-5+4c-1=0 3c+4c=5+1 7c=6 and c=6/7---Answer :D

OpenStudy (hba):

Now start solving your question ?

OpenStudy (anonymous):

to be honest im looking at the screen right now and i have no clue wat u just showed me lol

OpenStudy (hba):

Is your expression equal to 0 ?

OpenStudy (anonymous):

idk...... im thinking no

OpenStudy (hba):

Ok then, 7c+12-5c-9 Now you have to arrange them like, 7c-5c+12-9

OpenStudy (hba):

I hope you got it till here

OpenStudy (anonymous):

so u just do the oppisite?

OpenStudy (hba):

No i just rearranged them.

OpenStudy (hba):

The one with the c together.

OpenStudy (anonymous):

oh

OpenStudy (hba):

For Example, 3d+9+6-4d Now tell me how would you rearrange them :)

OpenStudy (anonymous):

3d-4d+9+6?

OpenStudy (hba):

You are Awesome :D

OpenStudy (anonymous):

i wish my teacher taught like this when she teaches its like she is speaking anotheer language

OpenStudy (hba):

Now lets go the original question, 7d + 12– 4d – 3 Rearrange them

OpenStudy (hba):

@-Lexy-

OpenStudy (anonymous):

7d-4d+12-3......?

OpenStudy (hba):

Awesome Now lets do some addition,subraction :)

OpenStudy (hba):

@-Lexy- Go on

OpenStudy (anonymous):

7d-4d=3d i think 12-3=9

OpenStudy (hba):

Ok so what is your final expression :

OpenStudy (hba):

@-Lexy-

OpenStudy (anonymous):

3d+9

OpenStudy (hba):

You are Awesome @-Lexy- And You Say You dont know maths :D

OpenStudy (hba):

A medal for you.

OpenStudy (anonymous):

i seriously didnt know that i just guessed but i got it right :o yay me!

OpenStudy (hba):

Now Lets Discuss the first question, Commutative Law, (When you add up) a+b=b+a (a+b)+c=a+(b+c) Distributive Law, (when you multiply) (a*b)*c=a*(b*c)

OpenStudy (hba):

So what will be the answer to the first question ? @-Lexy-

OpenStudy (anonymous):

Associative Property of Multiplication

OpenStudy (anonymous):

i think

OpenStudy (hba):

Yeah that is the answer :D As I said You are Awesome :P

OpenStudy (anonymous):

yay :D

OpenStudy (anonymous):

well im off to world history now